OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + y = 6 6x + 2y = 8 0 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system. I got =-12 but no y or x???

OpenStudy (anonymous):

3x + y = 6 . . 6x + 2y = 8 -2(3x + y = 6) . .6x + 2y = 8 -6x - 2y = -12 6x + 2y = 8 -------------------- 0 = -4

OpenStudy (anonymous):

if you multiply the left side of the first equation by 2, you get the second equation but if you multiply 6 by 2 you do not get 8 so there is no solution

OpenStudy (anonymous):

That's never true, so there is no x value as there is no solution to the system.